Grade 10Mechanics

A certain force exerted for 1.4 seconds on a moving body increases its speed from 2.8m/s to 5.6m/s the same force is later applied for 3 seconds,the speed of body changes by

Jayesh
7 Years agoGrade 10

1 Answer

Khimraj
7 Years ago
scceleration due to force = (change in velocity)/(time taken) = (5.6-2.8)/1.4 = 2 m/s2
change in velocity  v-u = at = 2*3 = 6m/s
Hope it clears..............................................